Two particles of masses 9m and 16m are placed at separation d.At what distance from 9m gravitational field is zero.

Grade:11

1 Answers

Krishan Pareek
6 years ago
let the distance from 9m is x where E=0.here at this point total E is zero or the gravitational field by both the masses must be equal.9m/x2 =(16m)/(d-x)29/x2 = 16/(d-x)29d2+9x2-18xd = 16x29d2-7x2-18xd=07x2-18xd-9d2=07x2-21xd+3xd- 9d2=0x=3d/7